The length of a fence is 4 feet less than twice the width. Find the dimensions if 32 feet of fence is used along three sides of it.
Answer by JulietG(1812)

L = 2W - 4 [The length of a fence is 4 feet less than twice the width.]
.
Not quite sure what the question is asking -- "if 32 feet of fence is used along three sides of it." Does it mean that 32' is part of the perimeter? We'll assume so.
.
If we know that 32' is 2L+w or 2W+l (because the perimeter is (2L + 2W), we can solve both ways.
Let's do 2W + L = 32 first.
Substitute the value of L from the first equation.
2W + (2W - 4) = 32
4W - 4 = 32
Add 4 to each side
4W = 36
Divide by 4
W = 9
If the width is 9, then L = 2(9)-4, or 18-4, or 14
W = 9 and L = 14. The perimeter would be (2*9) + (2*14), or 18 + 28 = 46
.
Let's solve it the second way.
2L + w = 32
Substitute the value of L from the first equation.
2(2W - 4) + W = 32
Distribute
4W - 8 + W = 32
5W - 8 = 32
Add 8 to each side
5W = 40
Divide by 5
W = 8
If W is 8, then L is (2*8)-4 or 16-4, or 12
Two lengths for the perimeter = 24 + two widths = 16. 16+24=40